OpenStudy (anonymous):

What is the pattern that utilizes every number within each arrangement. The third arrangement will have the same pattern as the first two arrangements: 9 3 1 11 , 8 , 6 7 , 14, 10 ? , 21, 8

OpenStudy (fifciol):

9+8=11+6 , 3+14=7+10 , 1+22=? +8 so ? =14

OpenStudy (fifciol):



OpenStudy (anonymous):

Awesome-- thanks so much for drawing out the relationship, too!

OpenStudy (anonymous):